if a girl is half of her mother,s age and the mother is five years young than the father and their combined number of years is 100,what is the age of the girl,the mother and father respectively
Answer by ankor@dixie-net.com(22740)

if a girl is half of her mother,s age and the mother is five years young than the father and their combined number of years is 100,
what is the age of the girl,the mother and father respectively
:
g + m + f = 100
:
"if a girl is half of her mother,s age"
g = .5m
:
" and the mother is five years young than the father "
m = f - 5
or
f = m + 5
:
Replace g and f in the 1s equation
.5m + m + (m+5) = 100
2.5m = 100 - 5
2.5m = 95
m = 95/2.5
m = 38 yrs is Mom's age
:
You can find the other ages